OpenStudy (anonymous):

arctan(tan(7pi/6)) what is the exact value

OpenStudy (mertsj):

\[\frac{7\pi}{6}\]

OpenStudy (anonymous):

I was wrong

OpenStudy (mertsj):

\[\tan \frac{7\pi}{6}=\frac{\sqrt{3}}{3}\]

OpenStudy (anonymous):

\[\tan ^{-1}(\tan7\times\frac{\pi}{6})\]

OpenStudy (mertsj):

\[\tan^{-1} (\frac{\sqrt{3}}{3})=\frac{\pi}{6}, \frac{7\pi}{6}\]

OpenStudy (anonymous):

\[=\frac{\pi}{6}= 30. degree\]

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw